OpenStudy (anonymous):

A porch has the shape of a semicircle with a diameter of 7 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate pi. I THINK IT IS 38, DOES ANYONE ELSE AGREE I WILL GIVE MEDAL IF EXPLAINED PROPERLY

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

I was right on the last one right??

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

on this one i am 99.9% percent sure it is 38

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

do u agree

OpenStudy (anonymous):

you see, it was really 38.465, but it says estimate, so it is 38

OpenStudy (anonymous):

is that rounded to the nearest square foot?

OpenStudy (anonymous):

@SolomonZelman @sleepyhead314 @pooja195 @PlanSlam @kirbykirby @KJ4UTS @kmachelle @preteenpony @LittleRedRiding @AwesomeSays3! @witchp @EclipsedStar @xX_GamerGirl_Xx @Zyi6

OpenStudy (anonymous):

@mathmath333 , how do i round to the nearest square foot

OpenStudy (anonymous):

and everyone else

OpenStudy (mathmath333):

38 is wrong answer

OpenStudy (anonymous):

oohh

OpenStudy (anonymous):

how, what is it then, and how is it that answer

OpenStudy (anonymous):

plz explain

OpenStudy (mathmath333):

area of semicircle=\(\Large \frac{1}{2}\times \pi\times r^2\)

OpenStudy (jackofalltradez):

You forgot to divide by 2

OpenStudy (mathmath333):



OpenStudy (anonymous):

I got 642456.4658, and I think that might be a little of XD

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

3.5 for semi circle

OpenStudy (anonymous):

so is it 1/2 times 3.14 times 3.5 2??

OpenStudy (anonymous):

the 2 is small

OpenStudy (anonymous):

but the half of what??

OpenStudy (anonymous):

@mathmath333 , might you plz help

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@BlakeCothran

OpenStudy (mathmath333):

calculate again \(\large =\frac{1}{2}\times 3.14\times \frac{7}{2}^2\) \(radius =\dfrac{diameter}{2}\)

OpenStudy (anonymous):

You're on the right track so far

OpenStudy (anonymous):

@mathmath333 , is'nt it 38.465/2

OpenStudy (mathmath333):

yes it is

OpenStudy (anonymous):

that is 19

OpenStudy (mathmath333):

yes correct

OpenStudy (anonymous):

but how do i round it to the nearest half square

OpenStudy (anonymous):

foot

OpenStudy (anonymous):

or is it already the answer, which is 19

OpenStudy (mathmath333):

include upto two decimal places

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the answer only alows me to have 3 spaces

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

you need to decimal spaces in order to round

OpenStudy (anonymous):

two*

OpenStudy (mathmath333):

ok then include only one digit after decimal

OpenStudy (anonymous):

if i do that then it is 19.

OpenStudy (anonymous):

when you round your two decimal spaces you can include one decimal

OpenStudy (anonymous):

it question will not let me input anymore, thats what my teacher said, only 3 digits

OpenStudy (anonymous):

ok, i am just gonna say 19

OpenStudy (mathmath333):